OpenStudy (anonymous):

Gaurav has a salary of Rs.21,600 per month. In the first month of the year, he spends 30% of his income on clothing, 20% of the remaining on entertainment and spend some percentage of the remaining on food. He saves 1/3 of what he spends. Find the percentage of spending on food as a % of saving.

OpenStudy (anonymous):

21600*30%=6480,15120*20=3024,6480+3024=9504,9504+x/3=

OpenStudy (callisto):

Let x be the amount he spent on food. 21600*30% + 21600*(1-30%)*20% + x + (1/3) * (21600*30% + 21600*(1-30%)*20% + x) = 21600. Then, you can find x.

Directrix (directrix):

Income: 21,600 Spent on Clothes: .30 * (21,600) = 6480 Remaining: 15,120 Spent on Entertainment: .20*6480 = 1,296 Remaining: 13,824 Spent on Food: x% --> (x/100)*13,824 = 138.24*x Remaining: 13,824 - 138.24*x to be saved ------------------------- 13,824 - 138.24*x (Savings) = 1/3 ( Expenditures) 13,824 - 138.24*x = 1/3( 6480 + 1,296 + 138.24*x ) x = 60.94 or .60% on Food 1/3 or .33% for savings Find the percentage of spending on food as a % of saving. ------------------------------------------------ Food % / Savings % = .60% / .33% = @fancy1 Something is not correct with my work.

Directrix (directrix):

@Callisto Do you see the first step where my work ran amok? Thanks.

OpenStudy (anonymous):

yes.

Directrix (directrix):

(A) 19% (B) 20% (C) 21% (D) 30%

OpenStudy (callisto):

Amount spent on entertainment should be .20*(15,120).

OpenStudy (anonymous):

x=26784

OpenStudy (anonymous):

6480+3024+(6480+3024+x)/3=21600 38010+x=64800 x=26784

OpenStudy (anonymous):

24%

OpenStudy (anonymous):

right ?

Directrix (directrix):

Income: 21,600 Spent on Clothes: .30 * (21,600) = 6480 Remaining: 15,120 Spent on Entertainment: .20*15,120 = 3,024 Remaining: 12,096 Spent on Food: x% --> (x/100)*12,096 = 120.96*x Remaining: 12,096 - 120.96*x to be saved ------------- (Savings) = 1/3 ( Expenditures) 12,096 - 120.96*x = 1/3*( 6480 + 3,024 + 120.96*x ) x = 55.36 or .554% for food Food/Savings = .554/.333 which is incorrect.

Directrix (directrix):

@Callisto Do you see my second mistake? And, did you get 24%?

Directrix (directrix):

@fancy1 I don't follow this logic: 6480+3024+(6480+3024+x)/3=21600 38010+x=64800 x=26784

OpenStudy (callisto):

I didn't get 24%, but I got 124%...

OpenStudy (anonymous):

hey 100-124=24%

OpenStudy (callisto):

24% of saving \(\ne\) 124% of saving

Directrix (directrix):

I think 24% is incorrect. I'm looking at 20% but there is no support that is correct. I don't know what has run amok.

OpenStudy (anonymous):

what is this @Callisto i cant understand

OpenStudy (callisto):

Suppose you have $100 as your saving. 24% of your saving is 100* (24/100) = 24 124% of your saving is 100 * (124/100) = 124

OpenStudy (callisto):

Are the options given in the question?

Directrix (directrix):

Not by the OP. I found options along with this same question on the Web. One of the options is None of these on one site. Regardless, there is something I am not seeing in the problem. So, I don't know what to think.

OpenStudy (callisto):

"x = 55.36 or .554% for food" This part is right. Food/salary = 0.31 Saving/salary = 0.25 Food/Savings = .31/.25

OpenStudy (anonymous):

but ur concept is right?.but it is 24%of 100 in 124 means 24 % extra

OpenStudy (anonymous):

got it @Callisto ,@directrix .the percentage of food is 31%

OpenStudy (anonymous):

i think the option is none of these

Directrix (directrix):

@fancy1 What options were you given with the problem?

OpenStudy (anonymous):

u already written

Directrix (directrix):

I think something is not right here: (Savings) = 1/3 ( Expenditures) --> 12,096 - 120.96*x <--- = 1/3*( 6480 + 3,024 + 120.96*x ) ----------------------------------- 12,096 - 120.96*x is the remaining Rs. after expenditures but is not the amount Gaurav will save.

OpenStudy (callisto):

If he didn't save what he didn't spend after expenditures, then we would have just got an expression in x?!

Directrix (directrix):

I agree with that.

OpenStudy (callisto):

\[\frac{12096-120.96x}{\frac{1}{3}(6480+3024+120.96x)}\]

Directrix (directrix):

Look at this person's work: o.0 let monthly income be = x on clothing =.3x on entertainment = .2(.7x) =.14x on food = (.16x)k saving =( .3x .14x .16xk)*1 solving we get k =.31.16 putting the value and we get ...... on the food =.31x ,and saving =.25x % on food / %on saving = .06x/.25x =24% is the correct answer http://tcyonline.minglebox.com/discuss/ans/5289/gauravs-salary-21600-per-monthfirst-month-year-he-spends-30-his-income-clothing-20-remaining-entertainment-some-perce

OpenStudy (callisto):

..... 0.06 is .31-.25

OpenStudy (callisto):

I know I can get 24% by doing (0.31-0.25)/0.25, but I don't understand why.

Directrix (directrix):

I don't follow that, and I have lost faith in my own work. I am going to rest my brain and then start from scratch. I'm still hung up here in my assumption that the savings is all the money left after expenditutes and is equal to 1/3 the expenditures. 12,096 - 120.96*x = 1/3*( 6480 + 3,024 + 120.96*x ) I suppose there is no reason to assume that he saves all of what he does not spend. lol

OpenStudy (callisto):

I suppose we can make that assumption.

OpenStudy (anonymous):

from answer can u get the answer

OpenStudy (anonymous):

like x+2=5,ie x=3,means 3+2=5,like that